6. Calculate the magnetic moment for the following complexes: a. (CrCl] [4% marks] b. [VC_04)2(NH3)21 [5% marks)
Answer #1

į Magnetic moment calculation (9.) [crcly J. Onidation state of a Onidation state of C%=> x-4=-1 x=-1+4=3 is a weak field lig